# Generator-Based Fuzzers with Type-Based Targeted Mutation

### v2

@@ 0 Type: Set of fuzzer-controlled inputs $I_{c}=V^{*}$
0 Type: Set of fuzzer-controlled inputs $I_{c}=(V,EI,\mathcal{T}^{*})^{*}$
@@ 0 Data: Set of uncovered code targets $\overline{\mathcal{C}}_{\mathcal{T}}\subseteq\mathcal{K}$
0 Data: Map of typing distance $\mathcal{D}$: $\mathcal{T}\rightarrow\mathbb{N}$
0 Data: Map of target typing distance $\Gamma:\mathcal{K}\rightarrow\mathcal{D}^{*}$
@@ 7                $i_{c}^{\prime}\leftarrow$ mutate $(i_{c})$
8                $i_{c}^{\prime},\mathcal{D}^{\prime}\leftarrow$ mutate $(i_{c},\overline{\mathcal{C}}_{\mathcal{K}},\mathcal{D})$
@@ 15                              $\mathcal{S}\leftarrow\mathcal{S}\cup\{i_{c}\}$
16                               $\mathcal{C}_{total}\leftarrow\mathcal{C}_{total}\cup$ coverage
17                               $\overline{\mathcal{C}}_{\mathcal{K}}\leftarrow\overline{\mathcal{C}}_{\mathcal{K}}-$ coverage
@@ 16                               $\mathcal{C}_{total}\leftarrow\mathcal{C}_{total}\cup$ coverage
Alg. 1 shows the main algorithm of GBF, with our extension highlighted in grey. The algorithm takes two inputs: an input generator $g$ and a program $p$. The fuzzer provides the fuzzer-controlled input from the set of FCI $I_{c}$ to the generator $g$ which returns a program input from the set of inputs $I_{p}$. Program $p$ is tested on input from $I_{p}$, then terminates with either a Failure or a Success result [^1] We assume that $p$ is always terminating. In the implementation a timeout is imposed for each test. With ${V}$, and $\mathcal{T}$ representing the set of random byte values and the set of all types, respectively, the GBF defines the following four data structures: (1) the set of fuzzer-controlled input $I_{c}$, which is defined as a sequence of random values, (2-3) the set of successful and failed input $\mathcal{S},\mathcal{E}$, and (4) The set of all code coverage $\mathcal{C}_{total}$.
@@
GBF proceeds as follows. Initially, the set of inputs is initialized with a single randomly-chosen FCI (line 15). Also, the failed inputs and the set of total coverages are set to the empty sets (line 15). Then, the algorithm proceeds in the fuzzing process until a specified time budget expires (lines 16-30). At each step, one input $i_{c}$ is picked up from $\mathcal{S}$, then the number of candidates to be generated from $i_{c}$ is determined using the numCandidates heuristic (line 18). In each iteration, the selected input is mutated to generate $i_{c}^{\prime}$ (line 19). Mutation means introducing random changes to random locations within $i_{c}$. Next, the generator is invoked on the mutated inputs (line 21), then GBF runs the program $p$ on the constructed program input $i_{p}$, and collects the result and the coverage information (line 22). If the program fails, the program input $i_{c}$ is added to the set of failed inputs (lines 23-24). Otherwise, if new coverage was achieved, GBF saves $i_{c}$ in the successful inputs and updates its total coverage (lines 26-29).
Our type-based targeted-mutation extension modifies the definition of the fuzzer-controlled input $I_{c}$ (line 5). That is, in our extension the fuzzer-controlled is defined by the product type of the randomly generated values $V$, execution indexes $EI$ defining the unique dynamic execution location within the input generators at the point of generating a specified random value, and sequence of types $T^{*}$ capturing the return value of each method within the callstack at the point of generating the random value. In addition to that, our extension introduces four additional data structures: (1) A set of all application code targets ($\mathcal{K}$), (2) the set of uncovered code targets ($\overline{\mathcal{C}}_{\mathcal{K}}$), (3) the map of typing distance ($\mathcal{D}$), and (4) the map of code targets with their typing distances ($\Gamma$).
Our algorithm starts by analyzing the program $p$. Our analysis collects code targets to populate the influencing types and together with their distances in $\Gamma$ (line 13). Using this map, our extended algorithm initializes the set of uncovered code targets and unifies their typing distances (line 14). The distance of a type indicates how far a type is from the code target. The closer the distance of a type the more likely it will be selected by the fuzzer for mutation. On the other hand, the unification of types aims to associate a single distance to each type. This is because the same type can have different distances for different code targets. Note that, mutation does not control which branch it will flip, so when choosing what type of object to mutate, we should choose based on all the branches that a type might be relevant to. This is why we unify type distances among all code targets. Our extension uses a different mutation operation by running the mutation operation over the updated fuzzer-controlled input $i_{c}$, the uncovered code targets $\overline{\mathcal{C}}_{\mathcal{K}}$ and the map of typing distances $\mathcal{D}$. The result of this operation is a mutated input ($i_{c}^{\prime}$), and an updated typing distance map $\mathcal{D^{\prime}}$ (line 20). Then, GBF proceeds with running the program on the newly constructed program input, then if new coverage is achieved then the new coverage is removed from the set uncovered code target if matched (line 29). We discuss each component in detail in the following sections (Sec.-3.1 - Sec. 3.2).
3.1 Static Phase
1. v17 = arrayload v10[v26]
2. v19 = invokevirtual File.toPath()Path v17
3. v21 = invokestatic YearThumbnails.yearTaken(Path)I v19
4. if(ne,v21,v3)
@@ 3.1 Static Phase
1. (Ljava/util/GregorianCalendar, 2)
2. (Ljava/nio/file/attribute/FileTime, 4)
3. (Ljava/nio/file/attribute/BasicFileAttributes, 5)
4. (Ljava/nio/file/Path, 6)
5. (Ljava/io/File, 8)
6. ([Ljava/io/File, 9)
@@
In this step, our extended algorithm analyzes the application code to find all code targets. The idea is to direct the fuzzing toward covering these code targets, in effect giving more precedence to application code coverage. In our analysis, we define code target to be either of the two sides of a branching condition within the application code. Then, for each code target, we compute types likely to influence their decision. An influencing type is any type that can flow to the operands of the code target.
In the general sense, our analysis computes a variation of def-use relation among variables. The analysis processes an IR derived from the bytecode. We use the Wala [17] static analysis framework to collect influencing types for this analysis, whose IR is in a static single-assignment form [18]. For example, Fig. 3 shows a simplified Wala’s IR representation of line 12 in Fig. 1(a), where the imageFiles is first loaded. This corresponds to loading elements of the array within the for loop at line 11. Then, at line 2 the image file is retrieved using toPath(), followed by invoking the application method yearTaken() to find the year of the photo (line 3), then finally branching on the resulting output (line 4). The analysis consists of three passes: a variable dependency pass, a type collection pass, and a type unification pass.
- Variable dependency pass: In this pass, we create a dependency graph between variables using a variation of the traditional def-use dataflow. Our extension gathers variable dependencies through a depth-first traversal of the application’s call graph. To capture the variable dependency when a method invocation is encountered, we add more edges to the dependency graph to capture the dependencies between the actual and formal parameters of the method. Similarly, if there is a return statement within the method, then we create a dependency between the returned variable in the callee and that of the caller. Also, for non-static methods, we capture the def-use dependency between the variables passed to the arguments and the reference variable on which the method invocation is done. Our analysis can identify recursive calls but visits the method only once at each encounter. Note that, for any invocation of a non-application method, i.e., a library call or a Java standard library call, we avoid analyzing their body. For example, the edges (use $\rightarrow$ def) in the dependency graph for this pass for Fig. 3 are:
@@
${\tt farg({yearTaken})}\rightarrow{\tt v19},{\tt v21}\rightarrow{\tt ret({yearTaken})}\}\cup{\tt E_{yearTaken}}$
Observe that v21 also depends on EyearTaken: the set of variable-dependency edges computed from analyzing the yearTaken() method. By contrast, the toPath() analysis is skipped since it is not identified to be defined within the application code (within the user’s outermost package name).
@@
- Type Collection Pass: In this pass, our extension walks all conditions and computes their influencing types. This is done by walking the variable dependency graph from the previous step for each operand within a branch condition. That is, we traverse the dependency graph backward, collecting encountered variable types and their respective distances. The distance to an influencing type indicates its proximity to the target branch, reflecting its significance to the mutation process. The intuition here is that closer influencing types probably have a larger effect on the branch’s decision. Moreover, the closer the influencing type is, the more likely it is to be fine-grained, which means, that the mutation can now happen over a smaller subsequence within the FCI. For example, Fig. 4 shows a simplified subset snippet of the influencing types collected for the branch at line 12 in Fig. 1(a). Each line is a pair of a type and a distance. Note that there might be multiple types at the same distance. In this case, we can see that the influencing type GregorianCalendar is much closer than the File type. This suggests that changing subsequences within the FCI related to the generation of the GregorianCalendar is more likely to make necessary changes to cover both arms of the intended branch. Finally, note that we ignore Java String types, primitive types, and their corresponding wrapper classes, such as Double and Float. These types are too common and do not provide sufficient distinction for targeting specific conditions.
@@
- Type Unification Pass: In this pass, we merge all influencing types with their distances for all code targets into a unified structure ($\mathcal{D}$). If influencing types are unique among all code targets then the unified structure will be the set of all types with their distance. However, if an influencing type appears in multiple code targets with different distances, then the type is added to $\mathcal{D}$ with its distance in the largest code target; i.e., the code target with the largest number of influencing types. The intuition is that the more influencing types the code target has, the more information and a finer distinction about the types exist in this set. For example, if $(t,d_{1})$ and $(t,d_{2})$ are found in two code targets $k_{1}$, and $k_{2}$, then the unified structure will contain $t,d_{3}$ such that $d_{3}=d_{1},\text{ if }|k_{1}|>|k_{2}|$, otherwise $d_{3}=d_{2}$, where $|k|$ denotes to the number of influencing types associated with a code target. Later, distances of various influencing types are updated during the dynamic fuzzing process depending on whether a static influencing type could be matched with a dynamic type. The change in distances allows the fuzzer to prioritize or de-prioritize influencing types depending on whether they were useful dynamically.
@@
- Creating the Constant String Lookup Table: We observed that using constant strings from the application code as string values when generating strings in the program input, can improve the fuzzing process. In the implementation, we employ static analysis to gather reachable string constants from the call graph’s entry point. Subsequently, during the fuzzing process, we enable the fuzzer to alternate between these constants and generating fresh strings when generating string values.
3.2 Fuzzing Phase
@@
In this step, we use information about influencing types collected from the static phase to make type-targeted mutations. Recall that there is no direct connection between the FCI and the program input. To enable type-directed mutation, one must first establish this connection between the bytes within the FCI and the types of objects they are used in generating. However, annotating the FCI with types is not enough to ensure that bytes after mutation will be used to generate the same object type. This is because mutations in the FCI are likely to lead to a different control flow within the generators. Thus if the FCI is consumed by the input generators only by the order the bytes they appear in, then there is no guarantee that the same byte within the mutated FCI will generate the same object type. One can only achieve that if we keep track of the dynamic locations where bytes within FCI are used. We do this by using execution indexing. Execution indexing (EI) [19, 20] of a program provides an ordered, unique representation of the dynamic points within the execution (as with an execution trace). Execution indexing technique has many applications such as program alignment [21], and detecting deadlocks and concurrency failures [22, 23]. The main idea of execution indexing is that every program point is associated with a vector of code points that identify the unique dynamic index of a particular execution point. We use EI to ensure that bytes can only be reused in the same dynamic locations. Next, we outline the process of annotating bytes within the FCI with types, creating their execution indices, and utilizing this information for type-targeted mutations.
3.2.1 Using the Execution Index to annotate the FCI with Types
@@
To annotate the FCI with types, we intercept invocations of requesting a random number, as in r.nextInt(). Then, we collect the type of objects that are currently being constructed. We do that by collecting the return type of each method invocation in the current callstack. For example, we intercept the execution of r.nextInt(31) in line 23 in Fig. 2(b). Let’s assume that the subsequence of the FCI for this random selection is the number 9. Then, we associate this value with the return types of all method invocations types in the current callstack. In this case, the value 9 will be annotated with {GregorianCalendar, File} types; i.e., (9,{GregorianCalendar, File}).
@@
As mentioned before, annotating bytes of the FCI with types is not sufficient for the targeted mutation. This is because without associating each byte within the FCI to their corresponding execution points, after mutation, the same byte can be reused in a different dynamic location, generating a different object type. For example, consider the case where initially generating a fileName required that the StrGenerator use the first 10 bytes of the FCI. Suppose a mutation occurred that resulted in having StrGenerator using only the first 5 bytes of the FCI. In that case, the remaining 5 bytes from the original FCI, while being annotated with String types, will be used within the FileGenerator to create an object of File type. This is a mismatch between the object types created by the same FCI’s subsequence from one test to another.
@@
To avoid reusing bytes in the same position in different FCIs, we associate an execution index with every byte of the FCI. We use an execution indexing implementation in Zest [24]. The main idea of the execution index is that every program point is associated with a vector of code points that identify the dynamic EI of a particular execution point. In our extension, we use a relaxed definition of execution indexing [14] implemented in Zest, where the uniqueness property among different execution indices is not guaranteed. More precisely, let $m_{1},m_{2},..,m_{n}$ be the call stack when the execution of an expression $p$ is about to happen. The execution index of $p$ (written as $ei(p)$) is defined as a vector of pairs $[l_{i},q_{i}]$ such that ($0<i\leq n$), where $l_{i}$ is a label of the expression within method $m_{i}$, and where $q_{i}$ is the number of times the statement was invoked during the current invocation of $m_{i}$.
@@
To illustrate, consider the input generator at Fig. 2(a) where a file is created and its creation date is constructed by invoking the CalGenerator at line 13 (Fig. 2(b)). If lCal is created during the first iteration of the while loop on line 9, then the execution index of the random choice of selecting a day of a month (line 23 in Fig. 2(b)); i.e., $ei(23)=[13,1,23,1]$. However, if lCal is created during the second iteration of the while loop then the execution index of the random choice of the day of a month would be $ei(23)=[13,2,23,1]$. This is because the generate(r,..) method at line 13 in this case has been invoked twice in the context of FileWithAttributesGenerator.
@@
Using this approach the FCI ($I_{c}$) would be represented as a sequence of triples of the form $I_{c}=(V,EI,\mathcal{T}^{*})^{*}$. For example, continuing our above example of creating the lCal object line 13. Our extended FCI for selecting the day of the month (line 23) will be the triple of the form: (9, ei(23),{GregorianCalendar, File})
3.2.2 Type-Based Mutation

### v3

@@ 0 Type: Set of fuzzer-controlled inputs $I_{c}=V^{*}$
0 Type: Set of fuzzer-controlled inputs $I_{c}=(V,EI,\mathcal{T}^{*})^{*}$  # replacing line 4
@@ 0 Data: Set of uncovered code targets $\overline{\mathcal{C}}_{\mathcal{T}}\subseteq\mathcal{K}$
0 Data: Map of type distance $\mathcal{D}$: $\mathcal{T}\rightarrow\mathbb{N}$
0 Data: Map of target type distance $\Gamma:\mathcal{K}\rightarrow\mathcal{D}^{*}$
@@ 7                $i_{c}^{\prime}\leftarrow$ mutate $(i_{c})$
8                $i_{c}^{\prime},\mathcal{D}^{\prime}\leftarrow$ mutate $(i_{c},\overline{\mathcal{C}}_{\mathcal{K}},\mathcal{D})$  # replacing line 19
@@ 15                              $\mathcal{S}\leftarrow\mathcal{S}\cup\{i_{c}\}$
16                               $\mathcal{C}_{total}\leftarrow\mathcal{C}_{total}\cup\text{coverage}$
17                               $\overline{\mathcal{C}}_{\mathcal{K}}\leftarrow\overline{\mathcal{C}}_{\mathcal{K}}\smallsetminus\text{coverage}$
@@ 16                               $\mathcal{C}_{total}\leftarrow\mathcal{C}_{total}\cup\text{coverage}$
Figure 5: Overall Process of Type-Based Targeted Mutation
Fig. 5 outlines the general process of our technique. First, we collect the influencing types for all branching conditions in a static process. Then, during fuzz testing (the dynamic phase), we annotate segments within the FCI with the types of objects they contribute to generating. Then, the targeted mutation process uses the influencing types information to mutate segments of the FCI with a matching dynamic type. In the process, coverage is collected, and influencing types are updated to reflect their selection priority. Additionally, we enable using of existing string constants, collected during the static phase, as string values for string generators, as they usually present some magic values within the program under test. In the rest of this section, we present the formal algorithm for the GBF extended with our changes for the type-targeting purposes, then we discuss in detail each process within the static and the dynamic phases.
III-A Algorithm
@@
Algorithm  1 shows the main algorithm of GBF with our extension highlighted in grey. We start by describing how GBF works, then we will introduce our extensions. In general, GBF’s algorithm takes two inputs: an input generator $g$, and a program $p$. A generator is a function that given a FCI $I_{c}$, it outputs a program input $I_{p}$. Program $p$ is tested on an input from $I_{p}$, then terminates with either a Failure or a Success result [^1] We assume that $p$ is always terminating. In the implementation a timeout is imposed for each test.. ${V}$ represents the set of random values, the set of fuzzer-chosen input $I_{c}$, which is defined as a sequence of random values. $\mathcal{S},\mathcal{E}$ represent the set of successful and failed inputs, respectively. And finally, $\mathcal{C}_{total}$ represents the set of all code coverage.
@@
GBF proceeds as follows. First, the set of successful inputs is initialized with a single randomly chosen FCI (line 15), while the failed inputs and the total coverage sets are initialized to the empty sets (line 15). Then, the fuzzing loop proceeds until a specified time budget expires (lines 16-30). At each step, an input $i_{c}$ is selected from $\mathcal{S}$, and the number of candidates to generate from $i_{c}$ is determined using the numCandidates heuristic (line 18). This function decides how many mutants will be produced from $i_{c}$. While the specifics of numCandidates are omitted, the number of mutants generally depends on how interesting $i_{c}$ is–specifically, how many new coverages are seen when executed. In each iteration, the selected input is mutated to generate $i_{c}^{\prime}$ (line 19). Mutation is done by introducing random changes to random locations within $i_{c}$. Next, the generator is then invoked on the mutated inputs (line 21) to create program inputs $i_{p}$, which the program executes (line 22). The fuzzer collects the result of the test and the coverage information (line 22). If the test fails, the program input $i_{c}$ is added to the set of failed inputs (lines 23-24). Otherwise, if new coverage was obtained, then GBF saves $i_{c}$ in the successful inputs and total coverage is updated (lines 26-29).
Our extension introduces $\mathcal{T}$, which represents the set of all possible types, and the set of execution indexes $EI$, which captures the unique dynamic location of a particular execution point. Thus, the fuzzer-chosen input $I_{c}$ (line 5) is defined as a tuple consisting of the randomly generated values $V$, the execution indexes $EI$ that specify the unique dynamic execution location within the input generators at the moment a specific random value is generated, and the sequence of types $T^{*}$, which records the return value of each method in the call stack at the point of generating the random value.
In addition to that, our extension introduces four additional data structures: (1) A set of all code targets ($\mathcal{K}$), (2) the set of uncovered code targets ($\overline{\mathcal{C}}_{\mathcal{K}}$), (3) the map of types to their distances ($\mathcal{D}$), and (4) the map of code targets to their typing distances ($\Gamma$). Note that, the distance of a type indicates how far a type is from flowing into the code target. The intuition is that, the closer the distance of a type the more likely it is affecting the branching decision, and thus it is more likely to be selected by the fuzzer for mutation.
Before fuzzing starts, we analyze the program $p$ to collect influencing types together with their distances in $\Gamma$ (line 13), which initializes the set of uncovered code targets. Next, since a type can appear multiple times during the analysis with varying distances to different code targets, we unify their typing typing distances (line 14) by associating a single distance to each type.
Our most significant is happening on line 20. Here, our extension uses a type-based mutation strategy. This process takes the updated fuzzer-chosen input $i_{c}$, the uncovered code targets $\overline{\mathcal{C}}_{\mathcal{K}}$ and the map of typing distances $\mathcal{D}$. The result of this operation is a mutated input ($i_{c}^{\prime}$), and an updated typing distance map $\mathcal{D^{\prime}}$ (line 20), which can de/prioritize types depending on whether a statically identified types was matched with a dynamic one. Finally, GBF proceeds by running the program on the newly constructed program input. If new coverage is achieved then the newly covered code target(s) is removed from the set uncovered code target (line 29). We discuss each component in detail in the following sections (III-B - III-C).
III-B Static Phase
1.v17 = arrayload v10[v26]
2.v19 = invokevirtual File.toPath()Path v17
3.v21 = invokestatic YearThumbnails.yearTaken(Path)I v19
4.if(ne,v21,v3)
@@ III-B Static Phase
1.(Ljava/util/GregorianCalendar, 2)
2.(Ljava/nio/file/attribute/FileTime, 4)
3.(Ljava/nio/file/attribute/BasicFileAttributes, 5)
4.(Ljava/nio/file/Path, 6)
5.(Ljava/io/File, 8)
6.([Ljava/io/File, 9)
@@
In this step, our extended algorithm analyzes the application code to find all code targets. Using this analysis we can direct the fuzzer to cover the code targets identified in this step. For each code target, we compute types likely to influence their decision. An influencing type is any type that can flow to the operands of the branching condition of the code target.
Our analysis computes a variation of def-use relation among variables. We use the Wala [15] static analysis framework to collect influencing types for this analysis, whose IR is in a static single-assignment form [16] representing the bytecode. For example,  Fig. 6 shows a simplified Wala’s IR representation of line 12 in  Fig. 1(a), where the imageFiles is first loaded. This corresponds to loading elements of the array within the for loop at line 11. At Fig. 6:L2 the image file is retrieved using toPath(), followed by invoking the application method yearTaken() to find the year of the photo (Fig. 6:L3), then finally branching on the resulting output (Fig. 6:L4). Note that our analysis is not field-sensitive, for example, although the image file has three date fields, our analysis only captures the GregorianCalendar as one of the influencing types. The analysis consists of three passes: a variable dependency pass, a type collection pass, and a type unification pass.
Variable dependency pass: In this pass, we use a flow-sensitive analysis where we create a dependency graph between variables using a variation of the traditional def-use dataflow. Our extension gathers variable dependencies through a depth-first traversal of the application’s call graph. To capture the variable dependency when a method invocation is encountered, we add more edges to the dependency graph to capture the dependencies between the actual and formal parameters of the method. Similarly, if there is a return statement within the method, then we create a dependency between the returned variable in the callee and that of the caller. Also, for non-static methods, we capture the def-use dependency between the variables passed to the arguments and the reference variable on which the method invocation is done. Our analysis can identify recursive calls but visits the method only once at each encounter. Note that, for any invocation of a non-application method, i.e., a library call or a Java standard library call, we avoid analyzing their body. For example, the edges (use $\rightarrow$ def) in the dependency graph for this pass for  Fig. 6 are:
@@
${\tt farg({yearTaken})}\rightarrow{\tt v19},{\tt v21}\rightarrow{\tt ret({yearTaken})}\}$
$\cup{\tt E_{yearTaken}}$
Here, the first line captures the dependencies between def-use variables. The second line captures the dependencies encountered due to the method invocation of yearTaken(), where farg(yearTaken), and ret(yearTaken) are functions that return the variable name of the formal argument as well as the variable name of the return variable of the same method. Observe that v21 also depends on EyearTaken: the set of variable-dependency edges computed from analyzing the yearTaken() method. By contrast, the analysis of the body of toPath() is skipped since it is not identified to be defined within the application code (within the user’s outermost package name).
@@
Type Collection Pass: In this pass, our extension walks all conditions of code targets and computes their influencing types. This is done by walking the variable dependency graph from the previous step for each operand within the code target branch condition. That is, we traverse the dependency graph backward, collecting encountered variable types and their respective distances. The distance to an influencing type indicates its proximity to the target branch, reflecting its significance to the mutation process. The intuition here is that closer influencing types probably have a larger effect on the branch’s decision. Moreover, the closer the influencing type is, the more likely it is to be fine-grained, which means, that the mutation can now happen over a smaller subsequence within the FCI. For example,  Fig. 7 shows a simplified subset snippet of the influencing types collected for the branch at line 12 in  Fig. 1(a). Each line is a pair of a type and a distance. Note that there might be multiple types at the same distance. In this case, we can see that the influencing type GregorianCalendar is much closer than the File type. This suggests that changing subsequences within the FCI related to the generation of the GregorianCalendar is more likely to make necessary changes to cover both arms of the intended branch. Finally, note that we ignore collecting Java String types, primitive types, and their corresponding wrapper classes, such as Double and Float. This is because these types are too common and do not provide sufficient distinction for targeting specific conditions.
@@
Type Unification Pass: In this pass, we merge all influencing types with their distances for all code targets into a unified structure ($\mathcal{D}$). If influencing types are unique among all code targets then the unified structure will be their union. However, if an influencing type appears in multiple code targets with different distances, then the type is added to $\mathcal{D}$ with its distance in the largest code target; i.e., the code target with the largest number of influencing types. The intuition is that the more influencing types the code target has, the more likely the code target can be covered if its corresponding influencing types are tweaked. For example, if $(t,d_{1})$ and $(t,d_{2})$ are found in two code targets $k_{1}$, and $k_{2}$, then the unified structure will contain ($t,d_{3}$) such that $d_{3}=d_{1},\text{ if }|k_{1}|>|k_{2}|$, otherwise $d_{3}=d_{2}$, where $|k|$ denotes to the number of influencing types associated with a code target. Later, distances of various influencing types are updated during the dynamic fuzzing process depending on whether a static influencing type could be matched with a dynamic type. The change in distances allows the fuzzer to prioritize or de-prioritize influencing types depending on whether they were useful dynamically.
@@
Creating Constant String Lookup Table In this step, we implement a simple analysis to identify magic strings, i.e., string that have specific meanings in the program under test. Our analysis gather all reachable string constants from the call graph’s entry point and within the application code. Later, we enable the fuzzer to alternate between selecting magic strings and generating fresh ones, when generating string values.
III-C Dynamic Fuzzing Phase
@@
In this step, we use information about influencing types collected from the static phase to make type-targeted mutations. Recall that there is no direct connection between the FCI and the program input. To enable type-directed mutation, one must first establish this connection between segments within the FCI and the types of objects they are used in generating. Annotating the FCI with types alone does not guarantee that mutated values will generate the same object type. This is because introduced mutations can alter control flow within generators, meaning that values may not be consumed in the same order, which might result in creating different objects than intended. To ensure consistency, we must associate with each data point in the FCI the dynamic location within the generators where it was used. We do this by using execution indexing. Execution indexing (EI) [17, 18] of a program provides an ordered, unique representation of the dynamic trace points within the execution. Execution indexing technique has many applications such as program alignment [19], and detecting deadlocks and concurrency failures [20, 21]. The main idea of execution indexing is that every program point is associated with a vector of code points that identify the unique dynamic index of a particular execution point. We use EI to ensure that values can only be reused in the same dynamic locations. Next, we outline the process of annotating values within the FCI with types, creating their execution indices, and utilizing this information for type-targeted mutations.
III-C1 Using Execution Index to annotate the FCI with Types
@@
To annotate the FCI with types, we intercept invocations requesting a random number, as in r.nextInt(). Then, we collect the type of objects that are currently being constructed. We do that by collecting the return type of methods within callstack. For example, we intercept the execution of r.nextInt(31) in line 23 in  Fig. 3. Let’s assume that the entry within the FCI for this random selection is the number 9. Then, we associate this value with the return types of all method invocations types in the current callstack. In this case, the value 9 will be annotated with {GregorianCalendar, File} types; i.e., (9,{GregorianCalendar, File}). As previously noted, annotating FCI values with types alone is insufficient for targeted mutation.
@@
In targeted mutation, the goal is to introduce localized changes to objects of a specific type while preserving the rest of the input. To prevent sequential processing of the FCI, which could disrupt this localization, we annotate each value with its unique dynamic location in the code by computing its execution index. For example, consider the case where initially generating a fileName required that the StrGenerator use the first 10 values of the FCI. Suppose a mutation occurred that resulted in having StrGenerator using only the first 5 values of the FCI. In that case, the remaining 5 values from the original FCI, while being annotated with String types, will be used within the FileGenerator to create an object of File type. This is a mismatch between the object types created by the same FCI’s subsequence from one test to another.
@@
To avoid problem we associate an execution index with every value of the FCI. The main idea of the execution index is that every program point is associated with a vector of code points that identify the dynamic EI of a particular execution point. In our extension, we use a relaxed definition of execution indexing [14] implemented in Zest, where the uniqueness property among different execution indices is not guaranteed. More precisely, let $m_{1},m_{2},..,m_{n}$ be the callstack when the execution of an expression $p$ is about to happen. The execution index of $ei(p)$ is defined as a vector of pairs $[l_{i},q_{i}]$ such that ($0<i\leq n$), where $l_{i}$ is the label of the statement within method $m_{i}$, and where $q_{i}$ is the number of times the statement was invoked during the current invocation of $m_{i}$.
@@
To illustrate, consider the input generator at  Fig. 3 where a file is created and its creation date is constructed by invoking the CalGenerator at line 13 ( Fig. 3). If lCal is created during the first iteration of the while loop on line 9, then the execution index of the random choice of selecting a day of a month (line 23 in  Fig. 3); i.e., $ei(23)=[13,1,23,1]$. However, if lCal is created during the second iteration of the while loop then the execution index of the random choice of the day of a month would be $ei(23)=[13,2,23,1]$. This is because the generate(r,..) method at line 13 in this case has been invoked twice in the context of FileWithAttributesGenerator.
@@
Using this approach the FCI ($i_{c}$) is represented as a sequence of triples of the form $i_{c}\in(V,EI,\mathcal{T}^{*})^{*}$. For example, continuing our above example of creating the lCal object (line 13). Our extended FCI for selecting the day of the month (line 23) will be the triple of the form: $(9,ei(23),\{GregorianCalendar,File\})$ Note that this technique is most effective when there is a separate generator method for every level of object that the generator constructs, a common coding practice.
III-C2 Type-Based Mutation